please answer the question correct

Using the following post-2018 tax cut tax rate schedule for Federal and Arkansas State corporate taxes, answer the following for XZY Corporation:

a. What is the: (i) Federal marginal tax rate; (ii) Arkansas state marginal tax rate; and (iii) combined Federal and Arkansas state marginal tax rate for an Arkansas corporation that will earn $40,000 in 2018.

b. What will the company’s after-tax earnings be after accounting for combined Federal and Arkansas state corporate taxes?

c. What will the effective tax rate be after accounting for combined Federal and Arkansas state corporate taxes?

d. XYZ Corporation has cumulative preferred stock but stopped paying preferred dividends more two years ago. If the company has surplus cash flows this year, and the Board wants to pay preferred and common stock dividends beginning this year without making up for missed dividends from prior years, can it do so?

e. XYZ Corporation goes through an initial public offering (IPO) and the stock will be traded on NASDAQ. Is NASDAQ an exchange or OTC market? Will XYZ Corporation IPO shares initially trade in the primary or secondary market?
Answer #1

For XYZ corporation

Earning $40,000, the federal marginal tax rates for 2018 is 15% respectively.

Arkansas Government tax would be 6%. Total tax bracket both federal+state(Arkansas) will be 15+6=21%

Marginal tax rate for XYZ corp. is 21%.

Therefore after-tax earnings=$40,000(1-0.21)= $31,600

Tax expense=40,000-31600=8400

Effective tax rate= 8400/40000=21% (Note marginal and effective calculated values might be same here as flat tax brackets are charged for corporations with lower income However if the income would have been more than 50,000 tax rate would have been $7500+15% for the additional amount from 50,000, then marginal and effective rates would have been different)

D) XYZ corporation can pay a dividend on common equity without making up for the missed dividend, however, there is a catch here. First, the Corporation needs to pay a Preferred dividend to its preferred shareholders and should also pay for their missed payments and then can pay dividends for common equity for this year.

In simple words, Corporation needs to pay up Preferred shareholders all their dividends for past two years then only they can pay a dividend to common equity shareholders. However no need to pay missed dividends for common equity shareholders.

E)National Association for Securities Dealers Automated Quote (NASDAQ) is a Hybrid market which is an Exchange. There are certain rules necessary to list itself on NASDAQ. If listed on NASDAQ the IPO shares will initially trade in a primary market.